OpenStudy (anonymous):

Whats the mass of 6.67 ml of acetone

OpenStudy (anonymous):

Or rather, how od i solve it? ( i wanna learn :P)

OpenStudy (anonymous):

ρ = m / v Keterangan : ρ = Massa jenis (kg/m3) m = massa (kg) v = volume (m3)

OpenStudy (anonymous):

i dunno masa jenis in english languange, you can find acetone \[\rho\]

OpenStudy (anonymous):

\[\rho acetone = 0.7925 g/ml\] mass = \[mass = \rho \times V\] = 0.7925 g/ml x 6.67 ml = 5.29 g convert gram into kilogram

OpenStudy (anonymous):

\[\rho \] means the mass density of substance

OpenStudy (anonymous):

:) got it. thank u

OpenStudy (anonymous):

youre welcome